# Parametric Operator Inference to Simulate the Purging Process in Semiconductor Manufacturing

### v1

@@ 3.1.4 ROM Interpolation
Once the reduced operators $\widehat{\mathbf{O}}_{1},\ldots,\widehat{\mathbf{O}}_{d}$ are learned for all $d$ parameters $\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{d}$, the reduced model for $\boldsymbol{\mu}\in\mathcal{M}$ is derived via elementwise interpolation [29] as
@@
For the implementation of elementwise interpolation between the parameter vectors $\boldsymbol{\mu}_{\ell}$ and the reduced operators $\widehat{\mathbf{O}}_{\ell}$, we use the LinearNDInterpolator module [30] in the $\mathtt{scipy}$ package [31] in Python. This interpolator constructs the ROM with reduced operators $\widehat{\mathbf{c}}^{*}$, $\widehat{\mathbf{A}}^{*}$, $\widehat{\mathbf{H}}^{*}$, and $\widehat{\mathbf{B}}^{*}$ for a given parameter vector $\boldsymbol{\mu}^{*}=(\mu_{q}^{*},\mu_{p}^{*})\in\mathcal{M}$, enabling state prediction for the new parameter unseen during the training. For the implementation of the parametric ROM learning and state predictions, we use the opinf Python package version 0.5.11 [32].
@@ 3.2 Data Transformation for Multiscale Data
The purging flow behavior in the PECVD chamber is multiphysics and multiscale, involving variables with different orders of magnitude. This poses a substantial challenge in model training, as variables with larger magnitudes can dominate those with smaller scales. In our case, the original variables, temperature and velocities, differ by up to four orders of magnitude. To mitigate numerical issues, we pre-process the training data by scaling all variables to comparable orders of magnitude. Specifically, we apply mean subtraction to the pressure and temperature fields, and scale all variables to the range of [$-1,1$]. This results in the scaled data matrix $\mathbf{X}\in\mathbb{R}^{1{,}246{,}310\times 1{,}800}$, which is used to construct the OpInf ROM.
@@ 3.3 OpInf ROM Learning via Regularization
In the model learning process (4), we use Tikhonov regularization to avoid overfitting, which arises from potential noise, under-resolution of the data, the truncated global POD modes that result in unresolved system dynamics, and mis-specification of the ROMs as fully quadratic [23]. Thus, the individual regression problem for each parameter $\boldsymbol{\mu}_{\ell}$ in (3) becomes
@@
where $\mathbf{\Lambda}=\textnormal{diag}(\lambda_{1},\lambda_{1}\mathbf{I}_{r},\lambda_{2}\mathbf{I}_{r(r+1)/2},\lambda_{3}\mathbf{I}_{m})\in\mathbb{R}^{d(r,m)\times d(r,m)}$ is the diagonal matrix used for regularization. Here, $\mathbf{I}_{r}$, $\mathbf{I}_{r(r+1)/2}$, and $\mathbf{I}_{m}$ are identity matrices with dimensions $r$, $r(r+1)/2$, and $m$, respectively. Note that the Kronecker product $\otimes$ introduces scaling differences between the entries of the quadratic operator $\widehat{\mathbf{H}}$, and those in the constant vector $\widehat{\mathbf{c}}$ and the linear operator $\widehat{\mathbf{A}}$, when subjected to a single regularization parameter. It is therefore best practice to use different regularization parameters $\lambda_{1}>0$, $\lambda_{2}>0$, and $\lambda_{3}>0$, such that $\lambda_{1}$ penalizes $\widehat{\mathbf{c}}$ and $\widehat{\mathbf{A}}$, $\lambda_{2}$ penalizes $\widehat{\mathbf{H}}$, and $\lambda_{3}$ penalizes $\widehat{\mathbf{B}}$ [23]. To determine the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ for regularizing the least-squares problem in (6), we generate a uniformly spaced parameter grid on a logarithmic scale, where ($\lambda_{1},\lambda_{2},\lambda_{3})\in[10^{-3},10^{3}]\times[10^{0}\times 10^{6}]\times[10^{-3}\times 10^{3}]$, with each dimension discretized into seven values. We find the regularization hyperparameters that minimize the relative state error on the training data. However, since the small magnitudes of the scaled data $\mathbf{X}$ can skew small errors, we compute the relative state error with the unscaled variables. To prevent variables with higher orders of magnitude from dominating those with smaller magnitudes in their original scales, we compute the relative error for each variable separately as
@@
where $\mathbf{S}^{.,i}\in^{n_{x}\times K}$ is the $i$th variable ($i=1$ for pressure, $i=2$ for temperature, $i=3$ for y-velocity, and $i=4$ for radial velocity). Note that since $v_{x}$ and $v_{z}$ are axisymmetric, we combine their effects on the total relative state error by computing a single radial velocity: $v_{r}=\sqrt{v_{x}^{2}+v_{z}^{2}}$. For ease of error comparison, we compute the average of the four relative errors, as $\frac{1}{4}\sum_{i=1}^{4}\mathcal{E}^{\text{rel},i}$. We then select the hyperparameters that minimize this error, which we found to be ($\lambda_{1},\lambda_{2},\lambda_{3}$) $=$ ($10^{-3},10^{2},10^{-3}$).
@@ 4.1 Selection of ROM Dimension
We determine the ROM dimension $r$ using the energy metric, which is defined by the singular values of the data matrix. To achieve this, we first compute the SVD of the scaled training data matrix $\mathbf{X}\in^{N\times dK}$, as described in (2). However, the deterministic SVD of $\mathbf{X}$ has a complexity of $O(N^{2}(dK)+(dK)^{3})$, making it impractical due to the computational load that grows quadratically in $N$. Thus, we compute a randomized SVD via the Python package Scikit-learn [33, 34].
Figure 4: (a) Normalized singular value decay. (b) the cumulative energy $\mathcal{E}_{\text{cum}}(r)$ and the residual energy that represents the projection error $\mathcal{E}_{\text{proj}}(r)$, with respect to the reduction dimension $r$. The vertical black line illustrates the chosen ROM dimension $r=5$.
From the randomized SVD of $\mathbf{X}$, the cumulative energy $\mathcal{E}_{\text{cum}}(r)$ is defined as the ratio of the sum of squared singular values up to rank $r$ to the total sum of squared singular values: $\mathcal{E}_{\text{cum}}(r)=\sum_{\eta=1}^{r}\sigma_{\eta}^{2}/\sum_{\eta=1}^{r_{t}}\sigma_{\eta}^{2}$. Here, $\sigma_{\eta}$ ($\eta=1,2,\dots,r_{t}$) are the singular values of the scaled data matrix $\mathbf{X}$, and $r_{t}$ is a target rank when computing randomized SVD. This relates the low-rank dimension $r$ to the projection error of the scaled data matrix $\mathbf{X}$ as follows:
@@
which quantifies the accuracy with which the linear basis $\mathbf{V}_{r}$ reconstructs the original data $\mathbf{X}$ through the linear injection of the projected data $\widehat{\mathbf{X}}=\mathbf{V}_{r}^{\top}\mathbf{X}\in^{r\times dK}$. Figure 4 shows the normalized singular value decay and the cumulative energy $\mathcal{E}_{\text{cum}}(r)$ related to the projection energy $\mathcal{E}_{\text{proj}}(r)$ in terms of the ROM dimension $r$. We choose $r=5$ since it reasonably captures the FOM dynamics, while achieving fast ROM simulations. The five modes retain a cumulative energy of $99.9981\%$, and a residual energy of $1.9056\times 10^{-5}$.
@@ 4.2 Prediction Results of Parametric OpInf ROM Interpolation
Figure 5(a) shows the selection of datasets and their corresponding parameter values used for training and testing. As discussed in Sec. 2.4, nine datasets, i.e., 36% of the entire data, are used for learning the ROM. Figure 5(b) shows the projection errors for all 25 ROMs when the global POD basis $\mathbf{V}_{r}$ with $r=5$ modes is used, based on nine selected training datasets. The projection error for each dataset is computed using (8). The results demonstrate that our basis $\mathbf{V}_{r}$ is well suited for both the training and testing data, with a maximum projection error of $0.74\%$ for the parameters ($\mu_{q},\mu_{p}$) = ($0.56,0.80$). Figure 5(c) shows the average relative state error for all 25 datasets computed using (7). Among the nine training datasets, the parameter combination ($\mu_{q},\mu_{p}$) = ($0.56,1.00$) shows the highest average state error of $3.91$%. For the 16 testing datasets, the lowest average error of $2.53$% occurs at ($\mu_{q},\mu_{p}$) = ($0.67,0.80$) and ($0.89,0.80$), while the highest error of $9.32$% is observed at ($\mu_{q},\mu_{p}$) = ($0.56,0.95$). Overall, the parametric OpInf ROM maintains good predictive accuracy, with errors remaining below 10% across all 25 parameter conditions.
@@
Figures 67, and 8 show the flow fields of both the FOM and ROM predictions over time, along with their corresponding pointwise errors, defined as
@@
where $j$ and $k$ are the spatial and temporal indices, respectively. Thus, the error $\Upsilon_{j,k}^{i}$ is the normalized absolute error of each unscaled variable over the entire spatiotemporal domain. All figures represent a vertical cross-section of the $xy$-vertical plane where $z=0$ [m], as shown in Figure 2(a). To effectively compare the differences in scale, the velocity components are normalized by the maximum absolute value between $v_{y}$ and $v_{r}$. The case presented corresponds to the parameters $(\mu_{p},\mu_{q})=(0.56,0.95)$, which yields the largest average relative state error (see Figure 5(c)).
@@
We omit the pressure results since they are nearly uniform and exhibit no significant gradients across the entire vertical cross-sectional area. However, we note that the ROM captures the FOM pressure behavior, with a maximum pointwise error of 0.429% at $t=1s$. Figure 6 shows that the ROM also predicts the temperature with high accuracy, producing a maximum error of $0.19\%$. Although this maximum error occurs specifically under the wafer heater, the ROM captures the temperature distribution across the entire spatiotemporal domain, especially given the small variation in its scale.
The velocity components, however, exhibit relatively high errors, as shown in Figures 7 and 8. In particular, the y-directional and radial velocities show maximum pointwise errors of $18.64\%$ and $20.01\%$, respectively. This case corresponds to the parameter combination $(\mu_{p},\mu_{q})=(0.56,0.95)$, which produces the highest average error of $9.32\%$ across pressure, temperature, y-velocity, and radial velocity, indicating that the velocity errors are the dominant contributors to this overall error. To analyze more in detail, the comparison between Figures 7(a) and 7(b) shows that the ROM accurately captures the overall direction of the y-velocity of the purging flow. However, the error contours in Figure 7(c) highlight the difference in magnitude between the FOM and ROM predictions, with a maximum error of 18.64% near the chamber outlet at the final time step ($t=1s$). Similarly, Figures 8(a) and 8(b) demonstrate that the ROM captures the overall pattern of the radial velocity field. However, Figure 8(c) shows that the errors increase toward the edge of the wafer surface at $t=1s$. From a practical perspective, accurate predictions of flow fields in the specific region that directly impacts particle contamination are especially important. During the PECVD process, after the plasma is deactivated, particles are carried by the purging flow. Since the level of contamination is determined by the amount of particles deposited on the wafer surface, the flow dynamics directly above the wafer surface are critical for particle contamination control. Thus, we define the monitor location as the $xz$-plane positioned 1 mm above the wafer surface (see Figure 9) to assess the prediction accuracy in the region of particle contamination.
@@
Figure 9: Monitor location: a quarter circular $xz$-plane with a 150 mm radius, positioned 1 mm above the wafer surface.
@@
We compare the y-velocity and radial velocity flow fields at the monitor location for the case with the largest error, $(\mu_{p},\mu_{q})=(0.56,0.95)$, in Figures 10 and 11. The velocities are normalized by the maximum absolute value between $v_{y}$ and $v_{r}$, highlighting that the radial velocity has a much larger magnitude than the y-velocity. As shown in Figure 10(a), the y-velocity is nearly uniform across the monitor surface, except at the edge, where it changes due to the acceleration of the purging flow. As shown in Figure 10(b), the ROM captures this behavior well, which is also partially observed in the area on top of the heater in Figure 7(c). Moreover, Figure 10(c) demonstrates that the ROM provides accurate y-velocity predictions across the entire surface, with a maximum error of 0.16%. In contrast, the radial velocity exhibits higher errors, as shown in Figure 11(c), especially near the wafer edge at $t=1s$. In fact, the velocity components of the purging flow play a crucial role in controlling particle contamination. The y-velocity directs particles toward the wafer surface, while the radial velocity transports them from the center toward the edge. Consequently, particles near the center must travel a longer distance to reach the edge, driven by the combined action of the radial and y-velocity. Moreover, the lower radial-to-y velocity ratio near the wafer center increases the risk of particle contamination and reduces the efficiency of particle removal. Thus, the low errors for both radial and y-velocity near the wafer center, as shown in Figures 10(c) and 11(c), suggest that the ROM accurately captures the key flow dynamics in the region most vulnerable to particle contamination.
@@
Figure 10: Contour plots of the y-velocity at the monitor location for the case with the largest error, $(\mu_{p},\mu_{q})=(0.56,0.95)$. Each column represents $t=0s$, $0.5s$, and $1s$. (a) FOM, (b) ROM predictions, (c) Pointwise normalized absolute error of y-velocity.
@@
Figure 11: Contour plots of the radial velocity at the monitor location for the case with the largest error, $(\mu_{p},\mu_{q})=(0.56,0.95)$. Each column represents $t=0s$, $0.5s$, and $1s$. (a) FOM, (b) ROM predictions, (c) Pointwise normalized absolute error of radial velocity.
@@
We measure the CPU time of the OpInf ROM to compare it with that of the FOM to evaluate computational efficiency. The OpInf ROM simulation consists of two stages: the offline stage, where the model is constructed, and the online stage, where predictions are made using the learned ROM. To assess the computational efficiency of each stage and identify potential computational bottlenecks, we measure the CPU time separately for the offline and online stages. To account for variations in ROM simulation time, we compute the average online cost over 20 simulations for all 25 ROMs. Table 1 compares the measured CPU time and shows the computational speedup of the ROM. To account for variations in ROM simulation time depending on various factors, we take the average of the online costs over 20 ROM simulations for 25 parameter combinations. The ROM construction in the offline stage takes 1491.27 minutes. Note that the model is constructed with nine parameter combinations (see  Figure 5(a)). However, the online stage, which predicts the flow for all 25 parameter combinations (including both training and testing cases), takes only 1.24 minutes on average over 20 simulations. This corresponds to about a 142-times speedup compared to the FOM computation time of 176.88 minutes. In other words, for many-query situations where we need to evaluate the ROM more than $1491.27/176.88\approx 8$ times, it pays off to construct a ROM instead of running the FOM.
Table 1: CPU time. The online speedup factor is computed by dividing the FOM simulation cost by the online ROM simulation cost.
FOM ROM offline ROM online Online speedup factor

### v2

@@ 3.1.4 ROM Interpolation
Once the reduced operators $\widehat{\mathbf{O}}_{1},\ldots,\widehat{\mathbf{O}}_{d}$ are learned for all $d$ parameters $\boldsymbol{\mu}_{1},\ldots,\boldsymbol{\mu}_{d}$, the reduced model for $\boldsymbol{\mu}\in\mathcal{M}$ is derived via elementwise interpolation [peherstorfer2020sampling] as
@@
For the implementation of elementwise interpolation between the parameter vectors $\boldsymbol{\mu}_{\ell}$ and the reduced operators $\widehat{\mathbf{O}}_{\ell}$, we use the LinearNDInterpolator module [10.1145/235815.235821] in the $\mathtt{scipy}$ package [virtanen2020scipy] in Python. This interpolator constructs the ROM with reduced operators $\widehat{\mathbf{c}}^{*}$, $\widehat{\mathbf{A}}^{*}$, $\widehat{\mathbf{H}}^{*}$, and $\widehat{\mathbf{B}}^{*}$ for a given parameter vector $\boldsymbol{\mu}^{*}=(\mu_{q}^{*},\mu_{p}^{*})\in\mathcal{M}$, enabling state prediction for the new parameter unseen during the training. For the implementation of the parametric ROM learning and state predictions, we use the opinf Python package version 0.5.11 [OpInfGitHub].
@@ 3.2 Data Transformation for Multiscale Data
The purging flow behavior in the PECVD chamber is multiphysics and multiscale, involving variables with different orders of magnitude. This poses a substantial challenge in model training, as variables with larger magnitudes can dominate those with smaller scales. In our case, the original variables, temperature and velocities, differ by up to four orders of magnitude. To mitigate numerical issues, we pre-process the training data by scaling all variables to comparable orders of magnitude. Specifically, we apply mean subtraction to the pressure and temperature fields, and scale all variables to the range of [1,1-1,1]. This results in the scaled data matrix $\mathbf{X}\in\mathbb{R}^{1{,}246{,}310\times 1{,}800}$, which is used to construct the OpInf ROM.
@@ 3.3 OpInf ROM Learning via Regularization
In the model learning process (4), we use Tikhonov regularization to avoid overfitting, which arises from potential noise, under-resolution of the data, the truncated global POD modes that result in unresolved system dynamics, and mis-specification of the ROMs as fully quadratic [McQuarrie_2021]. Thus, the individual regression problem for each parameter $\boldsymbol{\mu}_{\ell}$ in (3) becomes
@@
where $\mathbf{\Lambda}=\textnormal{diag}(\lambda_{1},\lambda_{1}\mathbf{I}_{r},\lambda_{2}\mathbf{I}_{r(r+1)/2},\lambda_{3}\mathbf{I}_{m})\in\mathbb{R}^{d(r,m)\times d(r,m)}$ is the diagonal matrix used for regularization. Here, $\mathbf{I}_{r}$, $\mathbf{I}_{r(r+1)/2}$, and $\mathbf{I}_{m}$ are identity matrices with dimensions $r$, $r(r+1)/2$, and $m$, respectively. Note that the Kronecker product $\otimes$ introduces scaling differences between the entries of the quadratic operator $\widehat{\mathbf{H}}$, and those in the constant vector $\widehat{\mathbf{c}}$ and the linear operator $\widehat{\mathbf{A}}$, when subjected to a single regularization parameter. It is therefore best practice to use different regularization parameters $\lambda_{1}>0$, $\lambda_{2}>0$, and $\lambda_{3}>0$, such that $\lambda_{1}$ penalizes $\widehat{\mathbf{c}}$ and $\widehat{\mathbf{A}}$, $\lambda_{2}$ penalizes $\widehat{\mathbf{H}}$, and $\lambda_{3}$ penalizes $\widehat{\mathbf{B}}$ [McQuarrie_2021]. To determine the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ for regularizing the least-squares problem in (6), we generate a uniformly spaced parameter grid on a logarithmic scale, where ($\lambda_{1},\lambda_{2},\lambda_{3})\in[10^{-3},10^{3}]\times[10^{0}\times 10^{6}]\times[10^{-3}\times 10^{3}]$, with each dimension discretized into seven values. We find the regularization hyperparameters that minimize the relative state error on the training data. However, since the small magnitudes of the scaled data $\mathbf{X}$ can skew small errors, we compute the relative state error with the unscaled variables. To prevent variables with higher orders of magnitude from dominating those with smaller magnitudes in their original scales, we compute the relative error for each variable separately as
@@
where $\mathbf{S}^{.,i}\in^{n_{x}\times K}$ is the $i$th variable ($i=1$ for pressure, $i=2$ for temperature, $i=3$ for y-velocity, and $i=4$ for radial velocity). Note that since $v_{x}$ and $v_{z}$ are axisymmetric, we combine their effects on the total relative state error by computing a single radial velocity: $v_{r}=\sqrt{v_{x}^{2}+v_{z}^{2}}$. For ease of error comparison, we compute the average of the four relative errors, as $\frac{1}{4}\sum_{i=1}^{4}\mathcal{E}^{\text{rel},i}$. The selected hyperparameters are ($\lambda_{1},\lambda_{2},\lambda_{3}$) == ($10^{-3},10^{2},10^{-3}$), which minimizes the training error across all $d$ ROMs during the model learning process.
@@ 4.1 Selection of ROM Dimension
Figure 4: (a) Normalized singular value decay, (b) the cumulative energy $\mathcal{E}_{\text{cum}}(r)$ and the residual energy that represents the projection error $\mathcal{E}_{\text{proj}}(r)$, with respect to the reduction dimension $r$. The vertical black line illustrates the chosen ROM dimension $r=5$.
We determine the ROM dimension $r$ using the energy metric, which is defined by the singular values of the data matrix. To achieve this, we first compute the SVD of the scaled training data matrix $\mathbf{X}\in^{N\times dK}$, as described in (2). However, the deterministic SVD of $\mathbf{X}$ has a complexity of $O(N^{2}(dK)+(dK)^{3})$, making it impractical due to the computational load that grows quadratically in $N$. Thus, we compute a randomized SVD via the python package Scikit-learn [halko2011finding, pedregosa2018scikitlearnmachinelearningpython]. From the randomized SVD of $\mathbf{X}$, the cumulative energy $\mathcal{E}_{\text{cum}}(r)$ is defined as the ratio of the sum of squared singular values up to rank $r$ to the total sum of squared singular values: $\mathcal{E}_{\text{cum}}(r)=\sum_{\eta=1}^{r}\sigma_{\eta}^{2}/\sum_{\eta=1}^{r_{t}}\sigma_{\eta}^{2}$. Here, $\sigma_{\eta}$ ($\eta=1,2,\dots,r_{t}$) are the singular values of the scaled data matrix $\mathbf{X}$, and $r_{t}$ is a target rank when computing randomized SVD. This relates the low-rank dimension $r$ to the projection error of the scaled data matrix $\mathbf{X}$ as follows:
@@
which quantifies the accuracy with which the linear basis $\mathbf{V}_{r}$ reconstructs the original data $\mathbf{X}$ through the linear injection of the projected data $\widehat{\mathbf{X}}=\mathbf{V}_{r}^{\top}\mathbf{X}\in^{r\times dK}$. Fig. 4 shows the normalized singular value decay and the cumulative energy $\mathcal{E}_{\text{cum}}(r)$ related to the projection energy $\mathcal{E}_{\text{proj}}(r)$ in terms of the ROM dimension $r$. The ROM dimension is set to $r=5$, which retains a cumulative energy of $99.9981\%$ (residual energy of $1.9056\times 10^{-5}$), providing an accurate low-dimensional representation of the FOM dynamics while substantially reducing computational cost. We note that while increasing $r$ beyond 5 further reduces the residual energy, it does not necessarily improve predictive accuracy in this type of nonlinear flow application. We tested ROMs with $r=10$ and $r=15$ for a single parameter case; neither improved the relative state error (7) compared to $r=5$.
@@ 4.2 Prediction Results of Parametric OpInf ROM Interpolation
Fig. 5(a) shows the selection of datasets and their corresponding parameter values used for training and testing. As discussed in Sec. 2.4, nine datasets, i.e., 36% of the entire data, are used for learning the ROM. The nine training cases are distributed across the full range of the two-dimensional parameter domain $\mu_{p}\in[0.80,1.00]$ and $\mu_{q}\in[0.56,1.00]$, whose convex hull forms a rectangle covering all 16 testing cases. This ensures that the surrogate model is evaluated strictly in the interpolation regime. Fig. 5(b) shows the projection errors for all 25 ROMs when the global POD basis $\mathbf{V}_{r}$ with $r=5$ modes is used, based on nine selected training datasets. The projection error for each dataset is computed using (8). The results demonstrate that our basis $\mathbf{V}_{r}$ is well suited for both the training and testing data, with a maximum projection error of $0.74\%$ for the parameters ($\mu_{q},\mu_{p}$) = ($0.56,0.80$). Fig. 5(c) shows the average relative state error for all 25 datasets computed using (7). Among the nine training datasets, the parameter combination ($\mu_{q},\mu_{p}$) = ($0.56,1.00$) shows the highest average state error of 3.913.91%. For the 16 testing datasets, the lowest average error of 2.532.53% occurs at ($\mu_{q},\mu_{p}$) = ($0.67,0.80$) and ($0.89,0.80$), while the highest error of 9.329.32% is observed at ($\mu_{q},\mu_{p}$) = ($0.56,0.95$). Overall, the parametric OpInf ROM maintains good predictive accuracy, with errors remaining below 10% across all 25 parameter conditions.
@@
Figs. 6, 8, and 9 show the flow fields of both the FOM and ROM predictions over time, along with their corresponding pointwise errors, defined as
@@
where $j$ and $k$ are the spatial and temporal indices, respectively. Thus, the error $\Upsilon_{j,k}^{i}$ is the normalized absolute error of each unscaled variable over the entire spatiotemporal domain. All figures represent a vertical cross-section of the $xy$-vertical plane where $z=0$ mm, as shown in Fig. 2(a). To effectively compare the differences in scale, the velocity components are normalized by the maximum absolute value between $v_{y}$ and $v_{r}$. The case presented corresponds to the parameters $(\mu_{p},\mu_{q})=(0.56,0.95)$, which yields the largest average relative state error (see Fig. 5(c)).
@@
We omit the pressure results since they are nearly uniform and exhibit no significant gradients across the entire vertical cross-sectional area. However, we note that the ROM captures the FOM pressure behavior, with a maximum pointwise error of 0.429% at $t=1s$. Fig. 6 shows that the ROM also predicts the temperature with high accuracy, producing a maximum error of $0.19\%$. Although this maximum error occurs specifically under the wafer heater, the ROM captures the temperature distribution across the entire spatiotemporal domain, especially given the small variation in its scale.
The velocity components, however, exhibit relatively high errors, as shown in Fig. 8 and 9. In particular, the y-directional and radial velocities show maximum pointwise errors of $18.64\%$ and $20.01\%$, respectively. This case corresponds to the parameter combination $(\mu_{p},\mu_{q})=(0.56,0.95)$, which produces the highest average error of $9.32\%$ across pressure, temperature, y-velocity, and radial velocity, indicating that the velocity errors are the dominant contributors to this overall error. To analyze more in detail, the comparison between Figs. 8(a) and 8(b) shows that the ROM accurately captures the overall direction of the y-velocity of the purging flow. However, the error contours in Fig. 8(c) highlight the difference in magnitude between the FOM and ROM predictions, with a maximum error of 18.64% near the chamber outlet at the final time step ($t=1s$). Similarly, Figs. 9(a) and 9(b) demonstrate that the ROM captures the overall pattern of the radial velocity field. However, Fig. 9(c) shows that the errors increase toward the edge of the wafer surface at $t=1$s. From a practical perspective, accurate predictions of flow fields in the specific region that directly impacts particle contamination are especially important. During the PECVD process, after the plasma is deactivated, particles are carried by the purging flow. Since the level of contamination is determined by the amount of particles deposited on the wafer surface, the flow dynamics directly above the wafer surface are critical for particle contamination control. Thus, the monitor plane is defined as the $xz$-plane at the height of 1mm above the wafer surface (see Fig. 7), as this region is most susceptible to particle contamination and therefore most critical for prediction accuracy.
Figure 7: Monitor location: a quarter circular xzxz-plane with a 150mm radius, positioned 1 mm above the wafer surface.
@@
We compare the y-velocity and radial velocity flow fields at the monitor location for the case with the largest error, $(\mu_{p},\mu_{q})=(0.56,0.95)$, in Fig. 11. The velocities are normalized by the maximum absolute value between $v_{y}$ and $v_{r}$, highlighting that the radial velocity has a much larger magnitude than the y-velocity. As shown in Fig. 10(a), the y-velocity is nearly uniform across the monitor surface, except at the edge, where it changes due to the acceleration of the purging flow. As shown in Fig. 10(b), the ROM captures this behavior well, which is also partially observed in the area on top of the heater in Fig. 8(c). Moreover, Fig. 10(c) demonstrates that the ROM provides accurate y-velocity predictions across the entire surface, with a maximum error of 0.16%. In contrast, the radial velocity exhibits higher errors, as shown in Fig. 11(c), especially near the wafer edge at $t=1$s. Taken together, the ROM captures the overall trend of radial velocity fields on the wafer surface well with a small number of POD modes ($r=5$), but it falls short in accurately predicting very localized features where the FOM exhibits sharp gradients near the edge. This is a common feature of global linear modal approximations. While incorporating—heuristically—more local mode shapes or nonlinear manifolds could potentially alleviate this issue, such extensions significantly increase the complexity of the OpInf problem [GEELEN2023115717]. For further discussions on both intrusive and nonintrusive localized ROM frameworks, see [amsallem2012nonlinear, geelen2022localized, buhr2020localized].
From a purging perspective, the two dominant velocity components of the purging flow serve distinct roles: the y-velocity vertically directs particles toward the wafer surface, while the radial velocity transports them laterally from the center toward the edge and ultimately to the exhaust outlet. Consequently, particles near the center must travel a longer distance to reach the edge, driven by the combined action of the radial and y-velocity, making this region inherently more susceptible to contamination. Near the wafer center, where $(x,z)=(0,0)[\textnormal{m}]$, the y-velocity reaches its highest magnitude while the radial velocity is at its minimum, yielding a low radial-to-y velocity ratio (see Fig. 11). This flow configuration promotes particle deposition onto the wafer surface, increasing contamination risk. Conversely, near the wafer edge, the radial velocity is dominant, producing a high radial-to-y velocity ratio that efficiently sweeps particles toward the exhaust outlet, reducing local contamination risk. The spatial ROM prediction errors are consistent with these contamination risk regions. Despite the maximum y-velocity error of 18.64% occuring near the outlet (bottom of the chamber), the error within the wafer center region remains below 0.16%, see Figs. 8(c) and 10(c). This confirms that the elevated y-velocity errors are spatially confined to the outlet vicinity, a region of negligible contamination risk. Similarly, while the radial velocity exhibits its highest error of 11.11% near the wafer edge, this region poses relatively lower contamination risk due to the dominant radial flow that promotes efficient particle purging. Moreover, as shown in Fig. 11(c), the radial velocity error decreases progressively toward the wafer center, the region of the highest contamination risk. This further confirms that the ROM maintains high predictive accuracy precisely in the area that is most critical to particle contamination.
@@
Figure 10: Y-velocity contour plots at the monitor location for the case with the largest error, $(\mu_{p},\mu_{q})=(0.56,0.95)$. Each column represents $t=0s$, $0.5s$, and $1s$. (a) FOM, (b) ROM predictions, (c) Pointwise normalized absolute error.
@@
Figure 11: Radial velocity contour plots at the monitor location for the case with the largest error, $(\mu_{p},\mu_{q})=(0.56,0.95)$. Each column represents $t=0s$, $0.5s$, and $1s$. (a) FOM, (b) ROM predictions, (c) Pointwise normalized absolute error.
@@
We measure the CPU time of the OpInf ROM to compare it with that of the FOM to evaluate computational efficiency. The OpInf ROM simulation consists of two stages: the offline stage, where the model is constructed, and the online stage, where predictions are made using the learned ROM. To assess the computational efficiency of each stage and identify potential computational bottlenecks, we measure the CPU time separately for the offline and online stages. To account for variations in ROM simulation time, we compute the average online cost over 20 simulations for all 25 ROMs. Both FOM and ROM computations are done using 20 CPU cores on a 13th-generation Intel Core i5-13600K processor (3.50 GHz). Table˜1 compares the measured CPU time and shows the computational speedup of the ROM. To account for variations in ROM simulation time depending on various factors, we take the average of the online costs over 20 ROM simulations for 25 parameter combinations. The ROM construction in the offline stage takes 1491.27 minutes. Note that the model is constructed with nine parameter combinations (see Fig. 5(a)). However, the online stage, which predicts the flow for all 25 parameter combinations (including both training and testing cases), takes only 1.24 minutes on average over 20 simulations. This corresponds to about a 142-times speedup compared to the FOM computation time of 176.88 minutes. In other words, for many-query situations where we need to evaluate the ROM more than $1491.27/176.88\approx 8$ times, it pays off to construct a ROM instead of running the FOM.
Table 1: CPU time comparison. The online speedup factor is computed by dividing the FOM simulation cost by the online ROM simulation cost.
FOM ROM ROM Online
offline online speedup factor